what is the resultant of 12N force and 9 N force pulling an object and forming an angle of 60 degree with each other

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
As we know
 
R^2 = 12^2 + 9^2 + 2 * 12 * 9 * cos 60
 
R^2 = 144 + 81 + 108 =333
 
R = sqrt(333) = 18.24 Newton